OpenStudy (anonymous):

I have the answer can somone please check me .... Find (dy/dx) x^3+y^3=3x

OpenStudy (anonymous):

\(y^3=3x-x^3\) \(y=\sqrt[3]{3x-x^3}\) \(y'=-\frac{x-3x^2}{3\sqrt[3]{(3x-x^3)^2}}\)

OpenStudy (anonymous):

um I am really confused that is not what I got

OpenStudy (anonymous):

it is just chain rule.

OpenStudy (anonymous):

you could write it also like this: \(y=(3x-x^3)^{\frac{1}{3}}\) and diferentiate it

OpenStudy (anonymous):

\[\frac{ 1-x^2 }{ y^2}\]Yes but look at the question again the answer that I got is

OpenStudy (anonymous):

no idea how you got that

OpenStudy (anonymous):

well you have to find\[\frac{ dy }{ dx }\] of the x^3+y^3=3x

OpenStudy (anonymous):

so first express y as function of x

OpenStudy (anonymous):

ok got that

OpenStudy (anonymous):

or use implicit differentiation

OpenStudy (anonymous):

what is the final answer that you get for the overall problem

OpenStudy (anonymous):

I already typed it in my 1º answer

OpenStudy (anonymous):

can you please retype what you think the answer is again please

OpenStudy (anonymous):

is the last line of my first answer

OpenStudy (anonymous):

ok but all my multiple choice answers I don't have that to choose from any of them

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

one is 3+3x^2+3y^2

OpenStudy (anonymous):

another is \[\frac{ 1-3x^2 }{ y^2 }\]

OpenStudy (anonymous):

ups I made a typo there

OpenStudy (anonymous):

where

OpenStudy (anonymous):

can you retype your answer please.

OpenStudy (anonymous):

in the numerator should be 3-3x^2

OpenStudy (anonymous):

ok well I still don't have that for any of my choices

OpenStudy (anonymous):

another answer is 3x^2+3y^2

OpenStudy (anonymous):

another one is \[\frac{ x^2 }{ y^2}\]

OpenStudy (anonymous):

another choice is \[\frac{ 3-x^2 }{ y^2}\]

OpenStudy (anonymous):

another chice is \[\frac{ -x^2 }{ y^2 }\]

OpenStudy (anonymous):

you could also write it as: \(\huge \frac{1-x^2}{y^2}\)

OpenStudy (anonymous):

another choice is 3-3x^2-3y^2

OpenStudy (anonymous):

ok thanks I gave that choice for my first answer but you said it was not right

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

its ok thanks for your help:)

OpenStudy (anonymous):

yw